To perform a 1:10 dilution, one would add _________microliters of sample to 2.7 milliliters of diluent

Show the equations used
Answer #1

Dilution factor, DF = 1/dilution

Here, dilution is 1/10, so DF will be 10.

DF = final volume (FV)/ initial volume (IV)

FV = IV + volume of diluent

Put in formula,

10 = IV + 2.7 / IV

10 IV = IV + 2.7

9 IV = 2.7

IV = 2.7/9 = 0.3

So, to dilute 10 times, add 0.3 ul sample to 2.7 ul diluent.

If you feel confused, than replace IV with x.

Please rate.